What is the molar mass of an unknown gas if a 25.0 g sample at 715 mmHg and 18.5 degrees celsius occupies 16.2 L
1
Expert's answer
2020-10-28T08:43:40-0400

pV=nRT. n=pV/RT = 715 x 16.2/62,4 x 291.5 = 0,636792454

Mr=m/n= 25/0,636792454=40